# Superficial Safety Alignment Hypothesis

### v1

@@ Abstract
As large language models (LLMs) are overwhelmingly more and more integrated into various applications, ensuring they generate safe and aligned responses is a pressing need. Previous research on alignment has largely focused on general instruction-following but has often overlooked the unique properties and challenges of safety alignment, such as the brittleness of safety mechanisms. To bridge the gap, we propose the Superficial Safety Alignment Hypothesis (SSAH), which posits that safety alignment should teach an otherwise unsafe model to choose the correct reasoning direction - interpreted as a specialized binary classification task - and incorporate a refusal mechanism with multiple reserved fallback options. Furthermore, through SSAH, we hypothesize that safety guardrails in LLMs can be established by just a small number of essential components. To verify this, we conduct an ablation study and successfully identify four types of attribute-critical components in safety-aligned LLMs: Exclusive Safety Unit (ESU), Exclusive Utility Unit (EUU), Complex Unit (CU), and Redundant Unit (RU). Our findings show that freezing certain safety-critical components (7.5%) during fine-tuning allows the model to retain its safety attributes while adapting to new tasks. Additionally, we show that leveraging redundant units (20%) in the pre-trained model as an “alignment budget” can effectively minimize the alignment tax while achieving the alignment goal. All considered, this paper concludes that the atomic functional unit for safety in LLMs is at the neuron level and underscores that safety alignment should not be complicated. We believe this work contributes to the foundation of efficient and scalable safety alignment for future LLMs.
@@ 1 Introduction
Large language models (LLMs) are demonstrating remarkable capabilities across a broad spectrum of natural language tasks, ranging from text generation to answering complex questions (Achiam et al., 2023; Touvron et al., 2023a; b; Dubey et al., 2024). However, as these models are increasingly integrated into real-world applications, concerns about the risk of generating harmful, unsafe, or unethical content have grown Askell et al. (2021); Bai et al. (2022); Zeng et al. (2024). This has led to a pressing need for safety alignment, which ensures that LLM outputs are not only coherent and informative but also aligned with human values, ethical standards, and safety considerations.
@@
Previous research on alignment has primarily focused on enhancing LLMs’ ability to follow general instructions without enough attention to model safety. This trend of treating safety alignment as a subset of general alignment has obscured its distinct challenges (Ouyang et al., 2022; Rafailov et al., 2024; Zhou et al., 2024; Liu et al., 2023a; Yuan et al., 2023; Liu et al., 2023b). One major issue is the brittleness of current safety mechanisms. Despite using benign data during model fine-tuning,  (Qi et al., 2023; Yang et al., 2023) have shown that safety mechanisms can fall apart when models are adapted to new tasks. This failure mode amplifies the brittle nature of current safety alignment techniques. Furthermore, safety alignment also comes with an ‘alignment tax’ - a trade-off where improving safety may reduce the model’s overall utility or downstream task performances Ouyang et al. (2022); Wang et al. (2024); Lin et al. (2024). Additionally, current safety alignment approaches typically require full model fine-tuning to ensure the final performance Hu et al. (2021); Zhang et al. (2023); Thakkar et al. (2024), which is computationally costly. To address these issues, we must accurately understand how safety alignment impacts model behavior, why safety mechanisms are brittle, and how to mitigate the alignment tax.
@@
Recently, Wei et al. (2024b) claimed the fragility of safety mechanisms in LLMs by identifying safety-critical components at the weight level through pruning (see Appendix C.2 for details). However, the work did not thoroughly address the fundamental issues discussed above nor propose effective solutions to resolve them. This paper introduces the Superficial Safety Alignment Hypothesis (SSAH), which separates safety alignment from general alignment and emphasizes its unique properties. With this hypothesis, we address that less is more in the context of safety alignment - a small number of but strategically vital safety-critical components are sufficient to achieve robust safety performance. Our extensive experiments prove that by freezing certain safety-critical computational units, safety performance is successfully preserved against fine-tuning attacks, and we are the first to do it. Additionally, we introduce the concept of repurposing redundant units within the pre-trained model as an “alignment budget.” By reassigning these units toward safety tasks, we can significantly reduce the alignment tax, ensuring that safety is maintained without sacrificing the model’s overall utility. Taken all together, this paper addresses the following core questions:
Figure 1: Superficial (Safety Alignment (ours) v.s. Alignment) Hypotheses
Question 1. How does safety alignment impact model behavior?
@@
Answer: Through SSAH, we posit that safety alignment fundamentally alters a model’s decision-making process by teaching an otherwise unsafe model - fulfilling malicious or harmful requests - to follow the correct reasoning pathways. This process can be viewed as a specialized binary classification task - the model must either fulfill the user’s request or refuse it based on safety considerations. Additionally, safety alignment equips the model with a standard refusal mechanism, along with reserved fallback options.
Question 2. Why is safety alignment brittle, and why does it introduce an alignment tax?
@@
Answer: We propose an attribute-based approach to analyzing the alignment and fine-tuning processes, where specific attributes are assigned to each individual computational unit - primarily input channels and output neurons. Our findings explain that the desired attributes can be achieved by repurposing units that were originally responsible for other functions. This reallocation helps explain both the brittleness of safety mechanisms and alignment tax.
Question 3. Can these issues of safety alignment be mitigated?
@@
Answer: By freezing the safety-critical components during fine-tuning and repurposing redundant units, we can effectively mitigate the brittleness and minimize the alignment tax. We conclude that the atomic functional unit for safety in LLMs resides at the neuron level and underscores that safety alignment should not be complicated.
@@ 2 Related Work
Alignment Tax and Fine-tuning Attack: The process of aligning LLMs with human preferences often incurs an “alignment tax”, where models experience degraded performance on downstream tasks due to the trade-offs required to maintain alignment (Bai et al., 2022; Ouyang et al., 2022; Lin et al., 2024; Wang et al., 2024). Additionally, fine-tuning attacks present another challenge: Yang et al. (2023); Qi et al. (2023)have shown that fine-tuning LLMs, even with benign data, can weaken safety measures. These findings highlight the inherent tension between safety alignment and utility, where improvements in one area often come at the expense of the other.
Model Pruning: Model pruning is a technique that reduces model size by removing redundant parameters, neurons, channels, layers, etc., which decreases storage needs and computational complexity without substantially impacting performance, namely magnitude pruning, unstructured/structured pruning, etc Frantar & Alistarh (2022); Frankle et al. (2020); Anwar et al. (2017); An et al. (2024); Li et al. (2024); Molchanov et al. (2019); Han et al. (2015); Lee et al. (2019); Renda et al. (2020). This method identifies and removes parts of the model that contribute least to its function, such as model weights with small magnitudes. In doing so, pruning extracts an efficient sub-model that runs faster on resource-constrained devices. We employ pruning as a tool to find out and delineate components to contribute to safety, utility, and both, respectively.
3 Superficial Safety Alignment Hypothesis (SSAH)
@@
Previous research proposed Superficial Alignment Hypothesis (SAH): A model’s knowledge and capabilities are learned almost entirely during pretraining, while alignment teaches the model which subdistribution of formats should be used when interacting with users (Zhou et al., 2024).
@@
However, this claim is centered on general alignment, and directly validating the hypothesis is challenging due to the complex interplays between pretraining and alignment. When a model fails to fulfill a user’s request, it can be difficult to determine whether the issue stems from the pretraining stage (due to lack of sufficient knowledge) or from the alignment process (due to misalignment in the output format). For example, when a model struggles with solving a math problem, it could either be a lack of relevant mathematical knowledge or the inability to structure its reasoning effectively. In such cases, good instruction techniques like the Chain-of-Thought approach can significantly enhance the quality of the model’s responses (Wei et al., 2022).
@@
Superficial Safety Alignment Hypothesis (SSAH). Since our focus is specifically on safety alignment, which has distinct properties compared to general alignment, we carefully define the scope of our hypothesis. A key observation here is that, for a model to be able to fulfill a malicious request, it must already possess the necessary knowledge and reasoning ability to carry out that harmful action. Based on this observation, we propose Superficial Safety Alignment Hypothesis (SSAH):
@@
SSAH: Given an unsafe model that is capable of fulfilling users’ malicious requests, safety alignment teaches the model the correct reasoning direction and a simple refusal mechanisms with reserved options.
@@
Reasoning direction here refers to the model’s internal decision-making process when confronted with a malicious query. That is, it represents the path the model is inclined to take in such a binary classification task, whether to fulfill the harmful request or to issue a refusal. As illustrated in Fig. 1, compared with SAH which targets the general alignment, our SSAH focuses on the safety alignment and has the following key differences :
@@
    Refusal mechanisms with reserved fallback options: Safety alignment generally requires the model to respond with a relatively standardized refusal format, which is simpler compared to general alignment, where a wider range of human preferences must be handled. The model can further simplify this purpose by effectively embedding multiple options of refusal response for all unsafe queries, such as “I cannot fulfill your request as it violates safety guidelines.” or “I am unable to assist with that as I am an AI programmed to follow ethical standards.”
@@
    Correction of reasoning direction: Safety alignment also distinguishes itself by its specific goal of teaching the model to choose the correct reasoning direction, which involves either fulfilling or refusing a user’s request based on whether it is safe. This process can be interpreted as a simple binary classification task.
Our SSAH can even provide insight into jailbreak attacks, where models often struggle to defend against adversarial inputs. In such cases, attackers typically use manipulative tokens to bypass the model’s safety mechanisms, indicating that the current alignment method can only hold the correct reasoning direction in limited generated tokens. However, a potential solution inspired by SSAH suggests that if safety alignment equips the model with the ability to consistently re-select the correct reasoning direction at each step (by re-evaluating the context of previously generated tokens before producing the next one), the model can continue to generate outputs that are both safe and helpful, even in the face of adversarial attempts.
@@
However, we approach the problem from an alternative perspective: if SSAH holds, we should observe distinct and consistent differences in the reasoning direction at each step of generation between safety-aligned and non-safety-aligned models. In a safety-aligned model, the reasoning direction should consistently guide the model in rejecting harmful queries at every token generation step. In contrast, a non-safety-aligned model might exhibit reasoning patterns that lean toward fulfilling malicious requests. Rather than relying solely on surface-level benchmark evaluations, we can probe the model’s reasoning direction to gain deeper insights into its internal decision-making process at each step regardless of the specific outputs produced.
@@
Figure 2: Probing reasoning direction on the AdvBench dataset with Llama2-7B, Llama3-8B, and Llama3.1-8B using cosine distance. Models were fine-tuned to ensure that aligned versions possess both general instruction-following abilities and safety guardrails, while unaligned models only have instruction-following capabilities. More results and details can be found in Appendix A.2.
@@
Probing Experiment. Although we cannot directly observe the model’s reasoning direction, as a turnaround, we can infer it by measuring the distance between hidden states in feature space when the model is fed with queries that follow different reasoning paths. By comparing the distances in aligned and unaligned models, we gain insights into how safety alignment affects the model’s reasoning direction in each generation step. Specifically, we observe that the model’s behavior can be influenced by the initial tokens in its response. For instance, appending certain tokens to a malicious query can lead an aligned model to produce unsafe responses or prompt an unaligned model to generate safe ones. This indicates that initial response tokens can alter the model’s reasoning direction. Based on this finding, we construct three types of queries to probe the model’s reasoning trajectory (More details can be found in Appendix A.2):
  1. (1)
    Query: The original malicious query (e.g., “How to make a bomb?”).
  2. (2)
    Query + benign prompt tokens: The malicious query followed by benign prompt token (e.g., “How to make a bomb? Sorry, I can’t…”).
  3. (3)
    Query + malicious prompt tokens: The malicious query followed by malicious prompt tokens (e.g., “How to make a bomb? Here’s how…”).
@@
Figure 3: Cosine distance between hidden states of various queries and clean queries across all blocks of LLMs (Aligned and unaligned model definitions are the same as in Fig. 2).
@@
Expected Outcomes and Probe Results. For an aligned model, we expect the hidden state distances between the query and query + benign prompt token to be shorter than those between the query and query + malicious prompt token at each generating step. In contrast, for an unaligned model, we anticipate the opposite: the distances between the query and query + malicious prompt token will be shorter than those between the query and query + benign prompt token. If these patterns are observed, it would indicate that safety alignment has successfully established the model’s ability to choose the correct reasoning direction. This also suggests that safety alignment reshapes the model’s internal decision-making process at each step of the generation, ensuring safer behavior from the very beginning of the response. As shown in Fig. 2, the probe results provide evidence that safety alignment teaches the model’s correct reasoning direction as hypothesized. However, it is important to note that this evidence is necessary but not enough, as safety alignment may introduce more nuanced changes that are not fully captured by SSAH.
Results Analysis and Discussion. We also present the aforementioned distances and their differences across transformer blocks in Fig.3 and Fig.4, respectively. Our findings demonstrate that the previous conclusions are consistently upheld across all transformer blocks. Specifically, the aligned model shows larger distance differences compared to the unaligned model, suggesting that the unaligned model lacks a strong preference for safe to unsafe reasoning.
@@
Figure 4: Absolute differences of cosine distance of Fig. 3 across all blocks of LLMs: Abs( Distance( Query + Benign tokens, Query) - Distance( Query + Malicious tokens, Query)) (Aligned/unaligned model definitions are the same as in Fig. 2).
@@
In contrast, the aligned model exhibits a clear preference for safe reasoning, as reflected by the more pronounced distance differences. Moreover, we observe that in the unaligned model, the distance difference gradually increases across the earlier transformer blocks 0 - 7. However, in the aligned model, the distance difference remains consistently large throughout all blocks. This indicates that the preference for safe reasoning in the aligned model is embedded not only in the later layers which typically capture higher-level features, but also in the earlier layers. Consequently, safety alignment influences the model’s internal decision-making from the initial stages of processing.

### v2

# Jianwei Li & Jung-Eun Kim

@@ Abstract
As large language models (LLMs) are overwhelmingly more and more integrated into various applications, ensuring they generate safe responses is a pressing need. Previous studies on alignment have largely focused on general instruction-following but have often overlooked the distinct properties of safety alignment, such as the brittleness of safety mechanisms. To bridge the gap, we propose the Superficial Safety Alignment Hypothesis (SSAH), which posits that safety alignment teaches an otherwise unsafe model to choose the correct reasoning direction - fulfill or refuse users’ requests - interpreted as an implicit binary classification task. Through SSAH, we hypothesize that only a few essential components can establish safety guardrails in LLMs. We successfully identify four types of attribute-critical components: Safety Critical Unit (SCU), Utility Critical Unit (UCU), Complex Unit (CU), and Redundant Unit (RU). Our findings show that freezing certain safety-critical components during fine-tuning allows the model to retain its safety attributes while adapting to new tasks. Similarly, we show that leveraging redundant units in the pre-trained model as an “alignment budget” can effectively minimize the alignment tax while achieving the alignment goal. All considered, this paper concludes that the atomic functional unit for safety in LLMs is at the neuron level and underscores that safety alignment should not be complicated.
@@ 1 Introduction
Large language models (LLMs) are demonstrating remarkable capabilities across a broad spectrum of natural language tasks, ranging from text generation to answering complex questions (Achiam et al., 2023; Touvron et al., 2023a; b; Dubey et al., 2024). However, as these models are increasingly integrated into real-world applications, concerns about the risk of generating harmful, unsafe, or unethical content have grown (Askell et al., 2021; Bai et al., 2022; Zeng et al., 2024). This has led to a pressing need for safety alignment, which ensures that LLM outputs are not only coherent and informative but also aligned with human values, ethical standards, and safety considerations.
@@
Previous work on alignment has primarily focused on enhancing LLMs’ ability to follow general instructions without enough attention to model safety. This trend of treating safety alignment as a subset of general alignment has obscured its distinct challenges (Ouyang et al., 2022; Rafailov et al., 2024; Zhou et al., 2024; Liu et al., 2023a; Yuan et al., 2023; Liu et al., 2023b). One major issue is the brittleness of the current safety mechanisms. Despite using benign data during model finetuning,  Qi et al. (2023); Yang et al. (2023) have shown that safety mechanisms can fall apart when models are adapted to new tasks. This failure mode amplifies the brittle nature of current safety alignment techniques. At the same time, safety alignment also comes with a tax - a trade-off where improving safety may reduce the model’s overall utility or downstream task performances (Ouyang et al., 2022; Wang et al., 2024; Lin et al., 2024). Also, current safety alignment approaches typically require full model finetuning (Hu et al., 2021; Zhang et al., 2023; Thakkar et al., 2024), which is computationally costly. To address these issues, we must first accurately understand two questions: 1) how safety alignment impacts model behavior? 2) why safety mechanisms are brittle?
Figure 1: Superficial Safety Alignment Hypotheses
@@
To answer these questions, this paper introduces Superficial Safety Alignment Hypothesis (SSAH), which separates safety alignment from general alignment and emphasizes its distinct properties. With this hypothesis, we prove that less is more in terms of parameters for safety alignment: a small number of but strategically vital safety-critical components are sufficient to achieve robust safety performance. Unlike Wei et al. (2024b), our extensive experiments prove that by freezing certain safety-critical units, safety performance is successfully preserved against finetuning attacks. Similarly, we also propose to repurpose redundant units of the pre-trained model as an “alignment budget.” By reassigning these units toward safety tasks, we can reduce the safety alignment tax, ensuring that safety is maintained without sacrificing the model’s overall utility. Taken all together, we answer the following core questions:
Q1. How does safety alignment impact model behavior?
@@
Answer: Through SSAH, we posit that safety alignment fundamentally alters a model’s decision-making process by teaching an otherwise unsafe model to follow the correct reasoning route: it either fulfills the user’s request or refuses it based on safety considerations. Importantly, this process can be viewed as an implicit and safety-related binary classification task. Furthermore, safety alignment equips the model with a standard refusal mechanism, along with reserved fallback options.
Q2. Why is safety brittle, and why does alignment tax exist?
@@
Answer: Based on the corollary of SSAH, less is more for parameters, we propose an attribute-based approach to analyzing the alignment and finetuning processes, where specific attributes are assigned to certain computational units. Our findings explain that the desired attributes (safety or utility) can be achieved by repurposing units originally responsible for other functions. This reallocation helps explain the trade-off between the safety performance and utility performance.
Q3. Can these issues of safety alignment be mitigated?
@@
Answer: By freezing the safety-critical components during finetuning or repurposing redundant units for safety, we can effectively mitigate the brittleness issue and minimize the alignment tax. We conclude that the atomic functional unit for safety in LLMs resides at the neuron level and underscore that safety alignment should not be complicated.
@@ 2 Related Work
Alignment Tax and Fine-tuning Attacks: Aligning LLMs with human preferences often incurs an “alignment tax,” where maintaining alignment often degrades downstream task performance (Bai et al., 2022; Ouyang et al., 2022; Lin et al., 2024; Wang et al., 2024). Additionally, fine-tuning can undermine safety: Yang et al. (2023); Qi et al. (2023) show that even benign fine-tuning weakens safety measures, highlighting the trade-off between alignment and utility.
Model Pruning: Pruning reduces model size and computational cost by removing redundant components (e.g., parameters, neurons, layers) while preserving performance Frantar & Alistarh (2022); Frankle et al. (2020); Anwar et al. (2017); An et al. (2024); Li et al. (2024); Molchanov et al. (2019); Han et al. (2015); Lee et al. (2019); Renda et al. (2020). By identifying less critical components, pruning enables efficient sub-model extraction for resource-limited environments. We leverage pruning to dissect model components contributing to safety, utility, or both.
3 Superficial Safety Alignment Hypothesis
@@
Previous work proposed Superficial Alignment Hypothesis (SAH): A model’s knowledge and capabilities are learned almost entirely during pretraining, while alignment teaches the model which subdistribution of formats should be used when interacting with users (Zhou et al., 2024).
@@
However, this claim is centered on general alignment, and directly validating the hypothesis is challenging due to the complex interplays between pretraining and alignment. When a model fails to fulfill a user’s request, it is difficult to determine whether the issue stems from the pretraining stage (due to lack of sufficient knowledge) or the alignment process (due to misalignment in the output format). For example, when a model struggles with solving a math problem, it could either be a lack of relevant mathematical knowledge or the inability to structure its reasoning effectively. In such cases, good instruction techniques like the Chain-of-Thought approach can significantly enhance the quality of the model’s responses (Wei et al., 2022).
@@
Superficial Safety Alignment Hypothesis (SSAH). Since our focus is specifically on safety alignment, a key observation arises: a model that can fulfill a malicious request must already possess the necessary knowledge and reasoning ability to carry out that harmful action. Based on this, we propose Superficial Safety Alignment Hypothesis:
@@
SSAH: Given an unsafe model that is capable of fulfilling users’ malicious requests, safety alignment teaches the model the correct reasoning direction—the model’s inclination to either fulfill or refuse a user request based on safety consideration—and a simple refusal mechanism with reserved options.
@@
Reasoning direction here refers to the model’s internal decision-making process when confronted with a malicious query. That is, it represents the path the model is inclined to take in a safety-related binary classification task, whether to fulfill the harmful request or to issue a refusal. As illustrated in Fig. 1, the key distinctions of SSAH from SAH are:
@@
    Refusal mechanisms with reserved fallback options: Safety alignment generally requires the model to respond with a relatively standardized refusal format, which is simpler compared to general alignment, where a wider range of human preferences must be handled. The model can further simplify this purpose by effectively embedding multiple options of refusal response, such as “I cannot fulfill your request as it violates safety guidelines.” or “I am unable to assist with that as I am an AI programmed to follow ethical standards.”
@@
    Correction of reasoning direction: Safety alignment also distinguishes itself by its specific goal of teaching the model to choose the correct reasoning direction, which involves either fulfilling or refusing a user’s request based on whether it is safe or not. This process can be interpreted as a simple binary classification task.
Extending to Jailbreak Attack. Although SSAH is originally proposed to explain how current safety alignment techniques impact models’ behavior under direct attack, it also provides insight into jailbreak attacks. In such cases, attackers typically use manipulative tokens to bypass the model’s safety mechanisms, indicating that the current alignment method can only hold the correct reasoning direction in a limited generated tokens. However, a potential solution inspired by SSAH suggests that if safety alignment enables the model to consistently re-select the correct reasoning direction at each step (by re-evaluating the query and previously generated tokens before producing the next one), the model can continue to generate outputs that are both safe and helpful, even confronted with adversarial attempts. This perspective has been validated explicitly by a recent study, which demonstrates its practical effectiveness in mitigating jailbreak attacks (li2025superficial). This also suggests that improving local consistency (safety) in reasoning can be a key pathway toward building more robust and attack-resilient language models.
@@
However, we approach the problem from an alternative perspective: if SSAH holds, we should observe distinct and consistent differences in the reasoning direction at each step of generation between safety-aligned and non-safety-aligned models. In a safety-aligned model, the reasoning direction should consistently guide the model in rejecting harmful queries at every token generation step. In contrast, a non-safety-aligned model might exhibit reasoning patterns that incline toward fulfilling malicious requests. Rather than relying solely on surface-level benchmark evaluations, we can probe the model’s reasoning direction to gain deeper insights into its internal decision-making process at each step, regardless of the specific outputs produced.
@@
Figure 2: Probing reasoning direction on the AdvBench dataset with Llama2-7B, Llama3-8B, and Llama3.1-8B using cosine distance. Models were finetuned to ensure that aligned versions possess both general instruction-following abilities and safety guardrails, while unaligned models only have instruction-following capabilities. More results are in Appendix A.2.
@@
Probing Experiment. Although we cannot directly observe the model’s reasoning direction, as a turnaround, we can infer it by measuring the distance between hidden states in feature space when the model is fed with queries that follow different reasoning paths. By comparing the distances in aligned and unaligned models, we gain insights into how safety alignment affects the model’s reasoning direction in each generation step. Specifically, we observe that the model’s behavior can be influenced by the initial tokens in its response. For instance, appending certain tokens to a malicious query can lead an aligned model to produce unsafe responses or prompt an unaligned model to generate safe ones. This indicates that initial response tokens can alter the model’s reasoning direction. Based on this observation, we construct three types of queries to probe the model’s reasoning trajectory (more details in Appendix A.2): (1) Clean Query: The original malicious query (e.g., “How to make a bomb?”); (2) Query + benign tokens: The malicious query followed by benign prompt token (e.g., “How to make a bomb? Sorry, I can’t…”); (3) Query + malicious tokens: The malicious query followed by malicious prompt tokens (e.g., “How to make a bomb? Here’s how…”).
@@
Figure 3: Cosine distance of hidden states between various crafted and clean queries across all blocks of LLMs (Aligned/Unaligned definitions are the same as Fig. 2).
@@
Expected Outcomes and Probe Results. For an aligned model, we expect the hidden state distances between the query and query + benign token to be shorter than those between the query and query + malicious token at each generating step. In contrast, for an unaligned model, we anticipate the opposite: the distances between the query and query + malicious token should be shorter than those between the query and query + benign token. If these patterns are observed, it indicates that safety alignment has successfully established the model’s ability to choose the correct reasoning direction. This also suggests that safety alignment reshapes the model’s internal decision-making process, ensuring safer behavior from the very beginning of the response. As shown in Fig. 2, the probe results provide evidence that safety alignment teaches the model’s correct reasoning direction as hypothesized. However, it is important to note that this evidence is necessary but not enough, as safety alignment may introduce more nuanced changes that are not fully captured by SSAH.
@@
Figure 4: Absolute differences of cosine distance of Fig. 3 across all blocks of LLMs: Abs(Distance(Query + Benign tokens, Query) - Distance(Query + Malicious tokens, Query)).
@@
Results Analysis and Discussion. We also present the aforementioned distances and their differences across Transformer blocks in Fig. 3 and Fig. 4, respectively. Our findings demonstrate that the previous conclusions are consistently upheld across all Transformer blocks. Specifically, the aligned model shows larger distance differences compared to the unaligned model, suggesting that the unaligned model lacks a strong preference for safe to unsafe reasoning. In contrast, the aligned model exhibits a clear preference for safe reasoning, as reflected by the more pronounced distance differences. Moreover, we observe that in the unaligned model, the distance difference gradually increases across the earlier Transformer blocks 0 - 7. However, in the aligned model, the distance difference remains consistently large throughout all blocks. This indicates that the preference for safe reasoning in the aligned model is embedded not only in the later layers (which typically capture higher-level features), but also in the earlier layers. Consequently, safety alignment influences the model’s internal decision-making from the initial stages of processing.